what is the compound interest on 2650$ in 5 years at 7 and half percent per anum
Answer by Theo(13342)

f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

in your problem.

f is what you want to find.
p is equal to 2650.
r is equal to 7.5% per year.
divide that by 100 to get an interest rate of .075 per year.
n is equal to 5 years.

formula becomes = 2650 * (1 + .075) ^ 5, resulting in f = 3804.417714.